3. (6pts) Find a general solution of the problem. 1- y = { cost, where (t > 0)
Answer #1

4-3 y-kcest y - 2 y = t cost Integrating factor - e at SA = e sent agenti Its Homplete integral is given by y c I-F) = (QCIF) dette =$+3 +3costatic 4+3 = Sint + c H18= t3 sint tot